# From Cool Demos to Production-Ready FMware: Core Challenges and a Technology Roadmap

### v1

@@ V-F Built-in Quality
Also, efforts like SPDX 3.0 Dataset profile [110] and Datasheets [111] aimed at enabling compliance for AI-powered software fail to capture the diverse types of data in FMware such as synthetic data, and user feedback data and their compliance requirements. This omission makes it challenging to ensure legal compliance in production-ready FMware.
Finally the immature QA process for prompts and agent actions contributes to the lack of built-in quality. Most systems lack built-in prompt validation mechanisms, relying on ad-hoc testing [112], which provides insufficient examples during in-context learning and leaves FMware prone to unpredictable behavior in complex tasks.
@@
Finally, for legal compliance, we propose an FMware Bill of Materials (FMwareBOM) by extending the SPDX 3.0 AI and dataset profiles [110]. The FMwareBOM would track all components and licenses, including synthetic data, RLHF data, and user feedback, addressing FMware’s unique complexities. A framework that automatically generates FMwareBOMs and uses formal verification techniques like SMT solvers would ensure provable compliance with legal and regulatory requirements. Integrating FMwareBOM into the development process is critical for overcoming compliance challenges and enabling the deployment of production-ready FMware.
VI Conclusion

### v2

@@ 5.6. Built-in Quality
Also, efforts like SPDX 3.0 Dataset profile (Unknown, 2023) and Datasheets (Gebru et al., 2021) aimed at enabling compliance for AI-powered software fail to capture the diverse types of data in FMware such as synthetic data, and user feedback data and their compliance requirements. This omission makes it challenging to ensure legal compliance in production-ready FMware.
Finally the immature QA process for prompts and agent actions contributes to the lack of built-in quality. Most systems lack built-in prompt validation mechanisms, relying on ad-hoc testing (Kuchnik et al., 2023), which provides insufficient examples during in-context learning and leaves FMware prone to unpredictable behavior in complex tasks.
@@
Finally, for legal compliance, we propose an FMware Bill of Materials (FMwareBOM) by extending the SPDX 3.0 AI and dataset profiles (Unknown, 2023). The FMwareBOM would track all components and licenses, including synthetic data, RLHF data, and user feedback, addressing FMware’s unique complexities. A framework that automatically generates FMwareBOMs and uses formal verification techniques like SMT solvers would ensure provable compliance with legal and regulatory requirements. Integrating FMwareBOM into the development process is critical for overcoming compliance challenges and enabling the deployment of production-ready FMware.
6. Conclusion